# Lifelong Sequential Knowledge Editing without Model Degradation

### v1

@@ 6 Conclusion
In this paper we present two major drawbacks of existing knowledge editing methods - (i) overfitting on edited facts and (ii) continuous and disproportionate norm-growth of edited matrix. We also present insights into the inner workings of locate-then-edit methods and show that they achieve edit success using the short-cut of norm-increase, which increases the importance of the output of edited matrices. We then present ENCORE, a faster and stronger knowledge editing method compared to its predecessors. ENCORE is able to perform 10,000 sequential knowledge edits without noticeable loss of downstream performance in about 40% of the time taken by other algorithms for Llama3-8B. While many recent works have shown limitations of knowledge editing methods at scale, we show that with a better understanding and appropriate regularizations, these methods can indeed be scaled. We hope this brings further research and excitement into the field of knowledge editing.
Impact Statement
This paper advances the field of knowledge editing in LLMs by enabling scalable, updates to a model’s stored knowledge without significant model degradation. Such techniques could be beneficial for multiple practical and societal uses, including quick updating of time-sensitive facts such as changing a model’s knowledge following natural disasters or newly published research.
However, as with any technology, there are potential risks with this method. The ability to efficiently overwrite and retain specific facts raises concerns about malicious editing or the injection of wrong information. Further research on robust governance and the use of these frameworks is still needed. Overall, we believe this research can lead to more responsive, up-to-date, and ethically aligned LLMs provided that careful attention is paid to responsible use.
@@ Appendix A Knowledge Editing Metrics
    Generation Entropy (GE): measures the fluency of the model’s text generation post-edit. GE is computed as the weighted average of bi-gram and tri-gram entropies in the text generated by the edited model. A lower GE indicates repetitive text generation, a common failure mode (Meng et al., 2022a; Gupta et al., 2024b).
@@
    Score (S): introduced by (Meng et al., 2022a), this composite metric combines edit success, generalization, and locality into a single score. It is calculated as the harmonic mean of the Efficacy Score (ES), Paraphrase Score (PS), and Neighborhood Score (NS).
@@ Appendix B Experimental Detail on Overfitting During Knowledge Editing
In each of the experiments we performed, we passed average numbers over 1000 edited facts with a batch size of 1. We use the CounterFact dataset (Meng et al., 2022a) for all of these experiments. We show the result in table 1. As we can see from the table, the unedited fact token probability is pretty low but once we run the edited fact the probability increases to almost 1 for some cases. MPES brings the probability of fact recall for edited facts down to a more natural value, which prevents the overfitting problem that we present in this paper.
Appendix C Experimental Details on Activation Norm Growth
@@
For each of the 30,000 examples, we calculated the proportion of the activation norm at each layer. We then plotted the mean and standard deviation of these proportions for both the edited and unedited models in figures 8 - 13. As shown in our results, the proportion of activation norms for the layers that were edited is significantly higher than their neighboring layers. In fact, some of the edited layers show the highest proportions overall. The edited layers are highlighted with red color on the x-axis.
C.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
We want to understand the effect of excessive increase in the norm of a vector in a sum of vectors. First, let’s start with an easy example where we have a summation of two vectors, $\mathbb{s}=\mathbb{a}+\mathbb{b}$ and then there is excessive increase in the norm of the first vector, that is $\mathbb{s}=k\mathbb{a}+\mathbb{b}$ where $k$ is some positive scalar. To evaluate the effect of this increase, we analyses the tendencies of the sum $\mathbb{s}$ as k increases.
We first want to understand the norm of $||\mathbb{s}||$. We have the following :
$||\mathbb{s}||^{2}=||k\mathbb{a}+\mathbb{b}||^{2}=k^{2}||\mathbb{a}||^{2}+2k\mathbb{a}\cdot\mathbb{b}+||\mathbb{b}||^{2}$
From this we can clearly see that as $k$ increases the first term quadratic in $k$ will dominate. This means that as $k\to\infty$, $||\mathbb{s}||^{2}\to k^{2}||\mathbb{a}||^{2}$, or $||\mathbb{s}||\to k||\mathbb{a}||$, which is the norm of the new vector. Thus, as the norm of one of the vectors in the summation increases, the norm of the summation tends to the norm of that vector with increasing norm.
Next, we look at the tendencies of the orientation of the summation as the norm of one vector increases. Let $\theta$ be the angle between $\mathbb{s}$ and $k\mathbb{a}$. Then, the cosine of the angle between the summation and the new vector $k\mathbb{a}$ is as follows (note that angle between $\mathbb{s}$ and $k\mathbb{a}$ is same as the angle between $\mathbb{s}$ and $\mathbb{a}$):
$\cos{\theta}=\frac{\mathbb{s}\cdot\mathbb{a}}{||\mathbb{s}||||\mathbb{a}||}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{||k\mathbb{a}+\mathbb{b}||||\mathbb{a}||}$
In the limit of $k\to\infty$, $||\mathbb{s}||\to k||\mathbb{a}||$ as shown above. Thus, the cosine expression in the limit can be written as:
$\underset{k\to\infty}{\cos{\theta}}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}=\frac{k\mathbb{a}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
$=1+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
@@ C.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
Finally, the same proof can be generalied to a summation of multiple vectors, where $\mathbb{b}$ represents the sum of the remaining vectors.
@@
(c) Average Norm Proportion for GPT2-XL using EMMET
(d) Average Norm Proportion for GPT2-XL using MEMIT
@@
(c) Average Norm Proportion for Llama2-7B using EMMET
(d) Average Norm Proportion for Llama2-7B using MEMIT
@@
(c) Average Norm Proportion for Llama3-8B using EMMET
(d) Average Norm Proportion for Llama3-8B using MEMIT
@@
(b) Average Norm Proportion for GPT2-XL using Norm Constraint
(c) Average Norm Proportion for GPT2-XL using ENCORE
Figure 11: Activation norm growth for GPT2-XL using Norm Constraint and ENCORE.
@@
(b) Average Norm Proportion for Llama2-7B using Norm Constraint
(c) Average Norm Proportion for Llama2-7B using ENCORE
Figure 12: Activation norm growth for Llama2-7B using Norm Constraint and ENCORE.
@@
(b) Average Norm Proportion for Llama3-8B using Norm Constraint
(c) Average Norm Proportion for Llama3-8B using ENCORE
Figure 13: Activation norm growth for Llama3-8B using Norm Constraint and ENCORE.
Appendix D Proof for ENCORE Objective
@@
Table 5: Editing performance for EMMET with Llama2-7B with norm constraint.
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
EMMET baseline Llama2-7B 93.85 87.32 58.07 76.281 579.79
EMMET best-editing Llama2-7B 94.24 87.2 54.36 74.12 566.08
EMMET best-downstream Llama2-7B 90.35 85.97 53.65 72.57 541.16
Appendix E Norm-Constraint with Other Objectives
While theoretically norm-constraint can be added to both EMMET and AlphaEdit objectives, it doesn’t work well with these objectives. AlphaEdit already has a norm constraint in its objective which is still not able to alleviate norm growth as shown in the main paper. Moreover, when norm-constraint is used with the EMMET objective, we see no appreciable improvements in downstream performance but incur a penalty in editing scores.
We believe that this is the nature of how norm-constraint works. When the norm-constraint is added, it penalizes change in weights from the original weight values. One way to prevent loss of downstream performance is to set the hyperparameter for norm-constraint very high, in which case we will never lose downstream performance. But in this scenario, we will also not be able to edit the model. In the other case, if we don’t restrict norm growth, the editor becomes free to increase the norm of edited matrices, which causes anomalous behaviors as discussed in the main paper. Thus, the norm constraint always comes at the trade-off of editability.
For the MEMIT objective, the norm-constraint helps both editing and downstream performance and the algorithm is best able to make use of this constraint. However, this is not the case for other algorithms. The theoretical reasons why this happens are beyond the scope of this work. But we provide evidence for this trade-off empirically for EMMET below.
Table 5 shows the editing performance of the EMMET baseline without any norm-constraint compared against two cases where norm-constraint is added to its objective. In one case, we present the best editing score that can be achieved when a non-zero norm-constrain hyperparameter is added to the EMMET objective ($\lambda_{n}=20$). In this case, we clearly see that even the best editing score with norm constraint is worse than the baseline EMMET. Figure 14(b) shows the downstream performance for this case, which is even worse than the baseline (Figure 14(a)).
Secondly, in Table 5 we also show the editing score for the hyperparameter that achieves the best downstream performance with non-zero norm-constrain hyperparameter in the EMMET objective ($\lambda_{n}=100$). We clearly see a significant drop from baseline editing performance across all metrics. The downstream performance of this case is shown in Figure 14(c), which is not much different from the baseline (Figure 14(a)). Thus, we show that the norm-constraint does not work well with the EMMET objective.
(a) EMMET baseline
(b) Downstream performance for best editing score with norm constraint.
(c) Downstream peformance for best the scenario where we get best downstream performance with norm constraint.
Figure 14: Comparison between the effect of adding norm-constrain the EMMET objective. We see no appreciable improvement in downstream performace but incur loss in editing performance.
Appendix F Editing Hyperparameters and Hardware Details
Here we present the hyperparameters that we find and used for the MPES, Norm constraint, and ENCORE. Tables 6, 7 and 8 show the hyperparameters for CounterFact dataset and tables 9, 10 and 11 show the hyperparameters for the zsRE dataset. Additionally, there is one more hyperparameter that requires tuning, as stopping immediately when the target token becomes the most probable may not always be optimal. We define this hyperparameter as the probability cutoff, which determines how many additional steps we take before stopping. Specifically, a cutoff of $+n$ means that we continue for $n$ more steps after the target token first becomes the most probable.
@@ Appendix F Editing Hyperparameters and Hardware Details
Method Model $\lambda_{p}$ Probability Cut Off
EMMET + MPES GPT2-XL 10,000 +1
Llama2-7B 15,000 +0
Llama3-8B 15,000 +0
AlphaEdit + MPEs GPT2-XL 20,000 +1
Llama2-7B 15,000 +0
Llama3-8B 15,000 +0
MEMIT + MPEs GPT2-XL 20,000 +2
Llama2-7B 15,000 +1
Llama3-8B 15,000 +2
@@
Method Model $\lambda_{p}$ $\lambda_{n}$
Norm Constraint GPT2-XL 20,000 10
Llama2-7B 15,000 10
Llama3-8B 15,000 20
Table 8: Hyperparameters for ENCORE on CouterFact dataset
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
ENCORE GPT2-XL 20,000 10 +3
Llama2-7B 15,000 10 +2
Llama3-8B 15,000 20 +1
@@
Method Model $\lambda_{p}$ Probability Cut Off
EMMET + MPES GPT2-XL 10,000 +1
Llama2-7B 15,000 +1
Llama3-8B 15,000 +2
AlphaEdit + MPEs GPT2-XL 20,000 +1
Llama2-7B 15,000 +1
Llama3-8B 15,000 +3
MEMIT + MPEs GPT2-XL 20,000 +5
Llama2-7B 15,000 +1
Llama3-8B 15,000 +4
Table 10: Hyperparameters for Norm Constraint on zsRE dataset
Method Model $\lambda_{p}$ $\lambda_{n}$
Norm Constraint GPT2-XL 20,000 40
Llama2-7B 15,000 10
Llama3-8B 15,000 20
Table 11: Hyperparameters for ENCORE on zsRE dataset
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
ENCORE GPT2-XL 20,000 40 +4
Llama2-7B 15,000 10 +2
Llama3-8B 15,000 10 +3
Appendix G Norm growth Result
In this section we present the observation of norm growth during editing for different methods (including Norm Constraint and ENCORE) with different models. Figures 15 and 16(c) show the result for different methods for GPT2-XL. Figures 17 and 18(c) show the result for Llama2-7B. Lastly figures 19 and 20(c) show the result for Llama3-8B.

### v2

@@ 7 Conclusion
In this paper, we show that existing knowledge editing methods require appropriate regularization when scaled sequentially to a large number of edits. Specifically, we diagnose and address two core failure modes in large-scale knowledge editing, (1) over-optimization of intermediate activations and (2) uncontrolled norm growth of edited matrices. To mitigate these, we introduce stage-specific regularization techniques in the form of “most-probable earlys stopping” (MPES) and Frobenius Norm-Constraint, which generalize to multiple knowledge editing methods. With the combination of both these regularization methods, we are able to make 10,000 sequential edits while consistently maintaining the downstream performance levels of the unedited model.
8 Limitation
Although our method significantly improves editing speed, scaling beyond 10,000 edits still remains challenging due to compute constraints. Performing 10,000 sequential edits on Llama3-8B already required a full day of GPU time on our hardware. This is because running these edits is combined with a lot of analysis, including downstream performance evaluation and measuring editing accuracy metrics on past edited facts. In contrast, baseline methods such as MEMIT combined with LTI were substantially slower, making larger-scale comparisons infeasible. Future work could explore more efficient optimization techniques or hardware acceleration to push beyond this limit.
Acknowledgements
This work was supported in part by the NVIDIA Academic Grant Program award.
@@ Appendix A Knowledge Editing Metrics
    Generation Entropy (GE): measures the fluency of the model’s text generation post-edit. GE is computed as the weighted average of bi-gram and tri-gram entropies in the text generated by the edited model. A lower GE indicates repetitive text generation, a common failure mode Meng et al. (2022a); Gupta et al. (2024b).
@@
    Score (S): introduced by Meng et al. (2022a), this composite metric combines edit success, generalization, and locality into a single score. It is calculated as the harmonic mean of the Efficacy Score (ES), Paraphrase Score (PS), and Neighborhood Score (NS).
@@ Appendix B Experimental Detail on Overfitting During Knowledge Editing
    Edited fact probability with LTI - Here we also evaluate the average probability with which a model recalls an edited fact. In this case we used LTI during editing the fact.
  4. 4.
@@
In each of the experiments we performed, we passed average numbers over 1000 edited facts with a batch size of 1. We use the CounterFact dataset Meng et al. (2022a) for all of these experiments. We show the result in table 1. As we can see from the table, the unedited fact token probability is pretty low but once we run the edited fact the probability increases to almost 1 for some cases. MPES brings the probability of fact recall for edited facts down to a more natural value, which prevents the overfitting problem that we present in this paper.
Appendix C GPT2-XL Result
In this section we show the result for GPT2-XL on CounterFact dataset in Tables 4 and 5 and also the downstream average downstream performance in Figure 6. As we can see that the result for GPT2-XL also improves the results as well.
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy MEMIT 94.04 79.91 57.90 74.22 517.37 MEMIT + LTI 82.03 70.51 56.29 67.97 508.35 MEMIT + MPES 92.17 77.01 60.38 72.26 523.57 PRUNE 61.05 58.05 50.00 55.96 579.69 PRUNE + LTI 80.06 69.77 56.07 67.18 542.41 PRUNE + MPES 76.91 59.11 65.96 66.54 563.57 RECT 51.40 49.83 52.17 51.12 409.42 RECT + LTI 47.51 47.37 52.74 49.08 185.96 RECT + MPES 51.99 49.93 54.29 52.01 523.63 AlphaEdit 88.58 70.33 56.04 69.20 580.27 AlphaEdit + LTI 74.20 59.90 54.86 61.98 596.47 AlphaEdit + MPES 95.52 82.08 60.03 76.32 565.44
Table 4: Knowledge editing performance for GPT2-XL on the CounterFact dataset for different algorithms in combination with MPES.
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy MEMIT 94.04 79.91 57.90 74.22 517.37 PRUNE 61.05 58.05 50.00 55.96 579.69 RECT 51.40 49.83 52.17 51.12 409.42 AlphaEdit 88.58 70.33 56.04 69.20 580.27 AlphaEdit + w/o NC 47.37 47.37 52.63 49.00 573.40 MEMIT + LTI + NC 84.15 73.12 54.91 68.54 554.47 MEMIT + NC 93.89 80.9 58.00 74.53 504.68 MEMIT + MPES + NC 93.99 79.79 59.98 75.29 517.44
Table 5: Knowledge editing performance for GPT2-XL on the CounterFact dataset for different algorithms in combination with MPES and NC.
(a) GPT2-XL downstream performance comparison with MPES
(b) GPT2-XL downstream performance comparison with MPES + NC
Figure 6: Average downstream performance for during sequential editing with compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
Appendix D Experimental Details on Activation Norm Growth
@@
For each of the 30,000 examples, we calculated the proportion of the activation norm at each layer. We then plotted the mean and standard deviation of these proportions for both the edited and unedited models in figures 7 - 12. As shown in our results, the proportion of activation norms for the layers that were edited is significantly higher than their neighboring layers. In fact, some of the edited layers show the highest proportions overall. The edited layers are highlighted with red color on the x-axis.
D.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
We want to understand the effect of excessive increase in the norm of a vector in a sum of vectors. First, let’s start with an easy example where we have a summation of two vectors, $\mathbf{s}=\mathbf{a}+\mathbf{b}$ and then there is excessive increase in the norm of the first vector, that is $\mathbf{s}=k\mathbf{a}+\mathbf{b}$ where $k$ is some positive scalar. To evaluate the effect of this increase, we analyses the tendencies of the sum $\mathbf{s}$ as $k$ increases.
We first want to understand the norm of $||\mathbf{s}||$. We have the following :
$||\mathbf{s}||^{2}=||k\mathbf{a}+\mathbf{b}||^{2}=k^{2}||\mathbf{a}||^{2}+2k\mathbf{a}\cdot\mathbf{b}+||\mathbf{b}||^{2}$
From this we can clearly see that as $k$ increases the first term quadratic in $k$ will dominate. This means that as $k\to\infty$, $||\mathbf{s}||^{2}\to k^{2}||\mathbf{a}||^{2}$, or $||\mathbf{s}||\to k||\mathbf{a}||$, which is the norm of the new vector. Thus, as the norm of one of the vectors in the summation increases, the norm of the summation tends to the norm of that vector with increasing norm.
Next, we look at the tendencies of the orientation of the summation as the norm of one vector increases. Let $\theta$ be the angle between $\mathbf{s}$ and $k\mathbf{a}$. Then, the cosine of the angle between the summation and the new vector $k\mathbf{a}$ is as follows (note that angle between $\mathbf{s}$ and $k\mathbf{a}$ is same as the angle between $\mathbf{s}$ and $\mathbf{a}$):
$\cos{\theta}=\frac{\mathbf{s}\cdot\mathbf{a}}{||\mathbf{s}||||\mathbf{a}||}=\frac{(k\mathbf{a}+\mathbf{b})\cdot\mathbf{a}}{||k\mathbf{a}+\mathbf{b}||||\mathbf{a}||}$
In the limit of $k\to\infty$, $||\mathbf{s}||\to k||\mathbf{a}||$ as shown above. Thus, the cosine expression in the limit can be written as:
$\underset{k\to\infty}{\cos{\theta}}=\frac{(k\mathbf{a}+\mathbf{b})\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}=\frac{k\mathbf{a}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}+\frac{\mathbf{b}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}$
$=1+\frac{\mathbf{b}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}$
@@ D.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
Finally, the same proof can be generalied to a summation of multiple vectors, where $\mathbf{b}$ represents the sum of the remaining vectors.
@@
(c) Average Norm Proportion for GPT2-XL using MEMIT
@@
(c) Average Norm Proportion for Llama2-7B using MEMIT
@@
(c) Average Norm Proportion for Llama3-8B using MEMIT
@@
(b) Average Norm Proportion for GPT2-XL using NC
(c) Average Norm Proportion for GPT2-XL using MPES + NC
Figure 10: Activation norm growth for GPT2-XL using NC and MPES + NC.
@@
(b) Average Norm Proportion for Llama2-7B using NC
(c) Average Norm Proportion for Llama2-7B using MPES + NC
Figure 11: Activation norm growth for Llama2-7B using NC and MPES + NC.
@@
(b) Average Norm Proportion for Llama3-8B using NC
(c) Average Norm Proportion for Llama3-8B using MPES + NC
Figure 12: Activation norm growth for Llama3-8B using NC and MPES + NC.
Appendix E Internal Computations within Decoder-Only LLMs
The internal computations happening within a transformer-based decoder-only LLM is shown below:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (7)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (9)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (10)
The intermediate hidden vector between each layer, $h^{l}$, is also sometimes referred to as the residual stream. LN1 represents the first LayerNorm (or equivalently RMSNorm for Llama models) that acts just before the attention module and LN2 represents the second LayerNorm just before the MLP module. Att represents the self-attention module in an LLMs whereas the action of a traditional MLP module is written in terms of the individual weight matrices ($W_{fc},W_{proj}$). As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model.
Appendix F Proof for MEMIT + MPES + NC Objective
@@
Appendix G Editing Hyperparameters and Hardware Details
Here we present the hyperparameters that we find and used for the MPES, Norm constraint, and MPES + Norm constraint. Tables 6, 7 and 8 show the hyperparameters for CounterFact dataset and tables 9,10 and 11 show the hyperparameters for the zsRE dataset. Additionally, there is one more hyperparameter that requires tuning, as stopping immediately when the target token becomes the most probable may not always be optimal. We define this hyperparameter as the probability cutoff, which determines how many additional steps we take before stopping. Specifically, a cutoff of $+n$ means that we continue for $n$ more steps after the target token first becomes the most probable.
@@ Appendix G Editing Hyperparameters and Hardware Details
Method Model $\lambda_{p}$ Probability Cut Off AlphaEdit + MPES GPT2-XL 20,000 +1 Llama2-7B 15,000 +0 Llama3-8B 15,000 +0 MEMIT + MPES GPT2-XL 20,000 +2 Llama2-7B 15,000 +1 Llama3-8B 15,000 +2
@@
Method Model $\lambda_{p}$ $\lambda_{n}$ Norm Constraint GPT2-XL 20,000 10 Llama2-7B 15,000 10 Llama3-8B 15,000 20
@@
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off MPES + NC GPT2-XL 20,000 10 +3 Llama2-7B 15,000 10 +2 Llama3-8B 15,000 20 +1
Table 8: Hyperparameters for MPES + Norm Constraint on CouterFact dataset
Method Model $\lambda_{p}$ $\lambda_{n}$ Norm Constraint GPT2-XL 20,000 40 Llama2-7B 15,000 10 Llama3-8B 15,000 20
Table 9: Hyperparameters for Norm Constraint on zsRE dataset
Method Model $\lambda_{p}$ Probability Cut Off AlphaEdit + MPEs GPT2-XL 20,000 +1 Llama2-7B 15,000 +1 Llama3-8B 15,000 +3 MEMIT + MPEs GPT2-XL 20,000 +5 Llama2-7B 15,000 +1 Llama3-8B 15,000 +4
@@
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off MPES + NC GPT2-XL 20,000 40 +4 Llama2-7B 15,000 10 +2 Llama3-8B 15,000 10 +3
Table 11: Hyperparameters for MEMIT+ Norm Constraint on zsRE dataset
Appendix H Norm growth Result
In this section we present the observation of norm growth during editing for different methods (including Norm Constraint and MPES + Norm Constraint) with different models. Figures 13 and 14 show the result for different methods for GPT2-XL. Figures 15 and 16 show the result for Llama2-7B. Lastly figures 17 and 18 show the result for Llama3-8B.